Probability

13. A particular bank is interested in estimating the difference between the average Mastercard debt at two of its branches. Using independent samples from two branches the 90% confidence interval for the difference between the means debt was calculated to be -S5.43 with 90% confidence that: H1 H2-$64.78. We could conclude from this A) branch 2 customers have more avera ge debt than customers from branch B) branch 1 customers have more average debt than branch 2 customers C) the minimum average amount owed by branch 1 customers is $5.43 D) there is no significant difference between the average debt at the two branches 14.A belief exists that there is a positive relationship between productivity growth and employment growth. To test this belief, a sample of twelve OECD countries data was used and a simple linear regression equation was found. The critical t value at the 1% level of significance would be: A) 3.169 B) 2.764 C) 3.106 D) 2.718 15. 17% of the employees in a company have managerial positions, and 28% of the employees have MBA degrees. Also, 60% of the managers have MBAs, which of the following is TRUE? A) The probability that a randomly selected employee is a manager and has B) The probability that a randomly selected employee is not a manager is C) Being a manager and having a MBA degree is independent. a MBA degree is 0.202. 0.45. If an employee having a MBA degree is selected, the probability that the D) employee is a manager is 0.364
Answer #1

Ans:

13)As,both limits are negative,so branch 2 customers have more average debt than customers from branch 1.

Option A is correct.

14)

n=12

df=n-2=12-2=10

critical t value(right tailed)=TINV(0.02,10)=2.764

Option B is correct.

15)

P(manages)=0.17

P(MBA)=0.28

P(MBA/managers)=0.6

P(MBA and manager)=0.6*0.17=0.102

P(manager/MBA)=P(manager and MBA)/P(MBA)=0.102/0.28=0.364

Option D is correct.